# Accelerating Ill-conditioned Hankel Matrix RecoveryHanQin Cai, Longxiu Huang, Xiliang Lu, and Juntao You

### v1

@@ 3 Theoretical results
We present the theoretical results of the proposed HSNLD in this section. We will first define some convenient notations and metrics for the presentation. Denote $\bm{L}_{\star}:={\bm{U}}_{\star}\bm{\Sigma}_{\star}^{1/2}$ and $\bm{R}_{\star}:={\bm{V}}_{\star}\bm{\Sigma}_{\star}^{1/2}$ where ${\bm{U}}_{\star}\bm{\Sigma}_{\star}{\bm{V}}_{\star}^{*}=\mathcal{G}\bm{z}_{\star}$ is the compact SVD of the underlying Hankel matrix. Define the distance between $({\bm{L}}_{k},{\bm{R}}_{k})$ and $(\bm{L}_{\star},\bm{R}_{\star})$ to be
(8) $d_{k}^{2}:=$ $\mathrm{dist}^{2}({\bm{L}}_{k},{\bm{R}}_{k};\bm{L}_{\star},\bm{R}_{\star})$
(9) ::=:= $\inf_{{\bm{Q}}\in\mathrm{GL}(r,\mathbb{C})}\big{\|}\left({\bm{L}}_{k}{\bm{Q}}-\bm{L}_{\star}\right)\bm{\Sigma}_{\star}^{1/2}\big{\|}_{\mathrm{F}}^{2}+\big{\|}\left({\bm{R}}_{k}{\bm{Q}}^{-*}-\bm{R}_{\star}\right)\bm{\Sigma}_{\star}^{1/2}\big{\|}_{\mathrm{F}}^{2}.$
Therein, $\mathrm{GL}(r,\mathbb{C})$ denotes the general linear group of degree $r$ over $\mathbb{C}$ and the invertible matrix ${\bm{Q}}$ is the best alignment between $({\bm{L}}_{k},{\bm{R}}_{k})$ and $(\bm{L}_{\star},\bm{R}_{\star})$. [33, Lemma 22] has shown the best alignment ${\bm{Q}}$ always exists if $d_{k}$ is sufficiently small. In that case, the infimum in (8) can be replaced by the minimum, and the standard recovery error $\left\lVert{\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}\right\lVert_{\mathrm{F}}$ can be bounded by $d_{k}$ as stated in the following proposition.
Proposition 3.1.
For any $\varepsilon_{0}\in(0,1)$, if provided $d_{k}<\varepsilon_{0}\sigma^{\star}_{r}$, then it holds that
@@
Proof 3.2.
The proof of this proposition is deferred to Section 5.2.
The following Theorem 3.3 guarantees that our initialization (7) will start HSNLD sufficiently close to the ground truth, under some mild conditions. Thus, it is sufficient to bound $d_{k}$ for the recovery guarantee. In addition, our initial guess also satisfies the weak incoherence condition, which will be used and inherited iteratively in the convergence analysis. For ease of presentation, we denote the basin of attraction as:
$\mathcal{E}(\delta,C):=\{({\bm{L}},{\bm{R}})~{}|~{}\mathrm{dist}({\bm{L}},{\bm{R}};\bm{L}_{\star},\bm{R}_{\star})\leq\delta~{}~{}\textnormal{and}~{}~{}\max\{\left\lVert{\bm{L}}{\bm{R}}^{*}\right\lVert_{2,\infty},\left\lVert{\bm{R}}{\bm{L}}^{*}\right\lVert_{2,\infty}\}\leq C\}.$
Theorem 3.3.
For any $\varepsilon_{0}\in(0,1)$, let $C=c\sqrt{c_{s}\mu r/n}\sigma^{\star}_{1}$ with $c\geq 1+\varepsilon_{0}$, and suppose Sections 1.1, 1.1 and 1.1 hold with $m\geq c_{0}\varepsilon_{0}^{-2}c_{s}\mu\kappa^{2}r^{2}\log n$ and $\alpha\leq\varepsilon_{0}/(50c_{s}\mu r^{1.5}\kappa)$, where $c_{0}$ is some universal constant. Then, the initialization step (7) satisfies
$({\bm{L}}_{0},{\bm{R}}_{0})\in\mathcal{E}(\varepsilon_{0}\sigma^{\star}_{r},C),$
with probability at least $1-2n^{-2}$. In the case of full observation, probability rises to $1$.
Proof 3.4.
The proof of this theorem is deferred to Section 5.3.
@@ 3.1 Full observation
We start the analysis with the simpler full observation case, i.e., $\Omega=[n]$. Theorem 3.5 presents the local linear convergence of HSNLD.
Theorem 3.5.
Suppose Sections 1.1 and 1.1 hold with $\alpha\lesssim\mathcal{O}(1/(c_{s}\mu r\kappa^{2}))$. Choose parameters $C=c\sqrt{c_{s}\mu r/n}\sigma^{\star}_{1}$ with $c\geq 1.01$, $\gamma_{k}\in[1+1/b_{0},2]$ with some fixed $b_{0}\geq 1$, and $\eta\in(0,1]$. Under full observation, for any $({\bm{L}}_{k},{\bm{R}}_{k})\in\mathcal{E}(0.01\sigma^{\star}_{r},C)$, it holds
@@
Proof 3.6.
The proof of this theorem is deferred to Section 5.4.1.
Combining Propositions 3.1, 3.3 and 3.5, we have the recovery guarantee, with a linear convergence, for the proposed HSNLD under full observation.
Corollary 3.7 (Recovery guarantee with full observation).
Suppose Sections 1.1 and 1.1 hold with $\alpha\lesssim\mathcal{O}(1/(c_{s}\mu r\kappa^{2}))$. Choose parameters $C=c\sqrt{c_{s}\mu r/n}\sigma^{\star}_{1}$ with $c\geq 1.01$, $\gamma_{k}\in[1+1/b_{0},2]$ with some fixed $b_{0}\geq 1$, and $\eta\in(0,1]$. Then, under full observation, in $K=\mathcal{O}(\log(1/\varepsilon))$ iterations, the outputs of Algorithm 1 satisfy
@@ 3.2 Partial observation
Now we will process the partial observation cases. One major challenge is the dependency among the entries on the antidiagonals of the Hankel matrix, which causes difficulty in uniformly bounding the subsampling error (see Lemma 5.26). Now, we present the local convergence under partial observation.
Theorem 3.8.
Suppose Sections 1.1, 1.1 and 1.1 hold with $m\gtrsim\mathcal{O}(c_{s}^{2}\mu^{2}r^{2}\kappa^{4}\log n)$ and $\alpha\lesssim\mathcal{O}(1/(c_{s}\mu r\kappa^{2}))$. Choose the parameters $C=c\sqrt{c_{s}\mu r/n}\sigma^{\star}_{1}$ with $c\geq 1.01$, $\gamma_{k}\in[1+1/b_{0},2]$ with some fixed $b_{0}\geq 1$, and $\eta\in(0,0.6]$. For any fixed $({\bm{L}}_{k},{\bm{R}}_{k})\in\mathcal{E}(0.01\sigma^{\star}_{r},C)$, with probability at least $1-8n^{-2}$ the following results hold
Proof 3.9.
The proof of this theorem is deferred to Section 5.4.2.
Combining Propositions 3.1, 3.3 and 3.8, we have the recovery guarantee, with a linear convergence, for the proposed HSNLD under partial observation.
Corollary 3.10 (Recovery guarantee with partial observation).
Suppose Sections 1.1, 1.1 and 1.1 hold with $m\gtrsim\mathcal{O}(c_{s}^{2}\mu^{2}r^{2}\kappa^{4}\log n)$ and $\alpha\lesssim\mathcal{O}(1/(c_{s}\mu r\kappa^{2}))$. Let the observation set $\Omega$ be resampled independently every iteration. Choose parameters $C=c\sqrt{c_{s}\mu r/n}\sigma^{\star}_{1}$ with $c\geq 1.01$, $\gamma_{k}\in[1+1/b_{0},2]$ with some fixed $b_{0}\geq 1$, and $\eta\in(0,0.6]$. Then, in $K=\mathcal{O}(\log(1/\varepsilon))$ iterations, the outputs of Algorithm 1 satisfy
$\left\lVert{\bm{L}}_{K}{\bm{R}}_{K}^{*}-\mathcal{G}\bm{z}_{\star}\right\lVert_{\mathrm{F}}\leq 0.02(1-0.6\eta)^{K}\sigma^{\star}_{r}\leq\varepsilon$
with probability at least $1-\mathcal{O}(n^{-1})$, provided $\log(1/\varepsilon)\ll n$.
Remark 3.11.
Compared to the state-of-the-art HSGD which runs $\mathcal{O}(\kappa\log(1/\varepsilon))$ iterations for a $\varepsilon$-optimizer, the convergence rate of HSNLD is independent of the condition number $\kappa$. This makes HSNLD a powerful tool for ill-conditioned Hankel applications, e.g., spectrally sparse signal recovery with inseparable frequencies [27].
@@ 4 Numerical experiments
In this section, we will empirically demonstrate the performance of the proposed HSNLD with synthetic and real datasets. We compare our algorithm with HSGD [4], SAP [37], and RobustEMaC [18]. Parameters are set according to the original paper and further hand-tuned for best performance when applicable. Particularly, both HSNLD and HSGD use the same outlier sparsification parameters: $\gamma_{k}=1.05+0.45\times 0.95^{k}$ as suggested in [4]. Moreover, we estimate the incoherence parameter $\mu$ via one-step Cadzow method [1] for all tested algorithms. The resampling is not implemented for HSNLD as it is just a proof technique and is not needed in practice. All results were conducted on Matlab with Intel i9-12950HX and 64GB RAM. The code is available online at https://github.com/caesarcai/HSNLD.
@@ 4.1 Synthetic dataset
We generate an ill-conditioned Hankal matrix in its vector form $\bm{z}\in\mathbb{C}^{n}$ with $z_{t}=\sqrt{\varsigma_{t}}\sum^{r}_{j=1}a_{j}e^{2\pi\imath f_{j}t},$ where $\imath$ denotes the imaginary unit, $\{f_{j}\}$ is a randomly selected set of normalized off-grid frequencies, and $\{a_{j}\}$ is a set of $r$ evenly spaced numbers over $[1/\kappa,1]$. Thus, $\mathcal{G}\bm{z}\in\mathbb{C}^{n_{1}\times n_{2}}$ is exactly rank-$r$ with condition number $\kappa$. Then, $m$ observations are sampled uniformly without replacement from $\bm{z}$, and therein $\alpha m$ entries are randomly chosen for corruption. We add complex outliers to the corrupted entries whose real and imaginary parts are drawn uniformly over the intervals $[-10\mathbb{E}(|\mathrm{Re}(z_{t})|),10\mathbb{E}(|\mathrm{Re}(z_{t})|)]$ and $[-10\mathbb{E}(|\mathrm{Im}(z_{t})|),10\mathbb{E}(|\mathrm{Im}(z_{t})|)]$.
Phase transition. We demonstrate the recoverability and robustness of HSNLD via empirical phase transition under various settings. When an algorithm stops at $\|\mathcal{G}\Pi_{\Omega}\bm{z}_{k}+\mathcal{G}\Pi_{\Omega}\bm{s}_{k}-\mathcal{G}\Pi_{\Omega}\bm{f}\|_{\mathrm{F}}/\|\mathcal{G}\Pi_{\Omega}\bm{f}\|_{\mathrm{F}}\leq 10^{-5}$, we considered it a successful recovery if $\|\mathcal{G}\bm{z}_{k}-\mathcal{G}\bm{z}_{\star}\|_{\mathrm{F}}/\|\mathcal{G}\bm{z}_{\star}\|_{\mathrm{F}}\leq 10^{-3}$. For every problem setting, each algorithm runs for $20$ trials and reports the result as a pixel on the phase transition, where a white pixel means all $20$ trials were recovered and a black pixel means all $20$ trials were failed. The phase transition results are reported in Figs. 2, 4 and 3 where one can see HSNLD delivers the best recoverability and robustness for ill-conditioned problems.
@@
Convergence performance. In Fig. 5, we plot the relative error v.s. runtime for the tested algorithms with different condition numbers. All problems have dimension $n=2^{16}-1$ and the results are averaged over 5 trials. When the condition number is perfect, i.e., $\kappa=1$, the HSNLD has tied convergence performance with HSGD. When the condition number grows, HSGD’s performance drops dramatically while HSNLD takes no impact—this matches our theoretical results. Generally speaking, SAP’s convergence performance is not impacted by the condition number either; however, the truncated SVD it applies every iteration may fail to function with vast condition number. RobustEMaC is not tested here because its sublinear convergence is too slow for the comparison.
@@
4.2 Nuclear magnetic resonance dataset
We test HSNLD with nuclear magnetic resonance (NMR) signal recovery problem, a classic application of low-rank Hankel recovery. The 1D NMR signal being tested has a length of $32,768$ with an approximate rank of $40$ and condition number $30$. The observation rate is set to be $30\%$ and $10\%$ outliers are added. All algorithms stop at $|\mathrm{err}_{k+1}-\mathrm{err}_{k}|/\mathrm{err}_{k}<10^{-4}$. The recovery results are reported in Figs. 6 and 1. One can see that all three tested algorithms recovered the signal successfully while HSNLD is significantly faster.

### v2

@@ 3 Theoretical results
We present the theoretical results of the proposed HSNLD in this section. We will first define some convenient notations and metrics for the presentation. Denote $\bm{L}_{\star}:={\bm{U}}_{\star}\bm{\Sigma}_{\star}^{1/2}$ and $\bm{R}_{\star}:={\bm{V}}_{\star}\bm{\Sigma}_{\star}^{1/2}$ where ${\bm{U}}_{\star}\bm{\Sigma}_{\star}{\bm{V}}_{\star}^{*}=\mathcal{G}\bm{z}_{\star}$ is the compact SVD of the underlying Hankel matrix. Let $({\bm{L}}_{k},{\bm{R}}_{k})$ denote the update at the $k$-th iteration in Algorithm 1. Define the distance between $({\bm{L}}_{k},{\bm{R}}_{k})$ and $(\bm{L}_{\star},\bm{R}_{\star})$ to be
$d_{k}^{2}:=\mathrm{dist}^{2}({\bm{L}}_{k},{\bm{R}}_{k};\bm{L}_{\star},\bm{R}_{\star}):=\inf_{{\bm{Q}}\in\mathrm{GL}(r,\mathbb{C})}\big{\|}\left({\bm{L}}_{k}{\bm{Q}}-\bm{L}_{\star}\right)\bm{\Sigma}_{\star}^{1/2}\big{\|}_{\mathrm{F}}^{2}+\big{\|}\left({\bm{R}}_{k}{\bm{Q}}^{-*}-\bm{R}_{\star}\right)\bm{\Sigma}_{\star}^{1/2}\big{\|}_{\mathrm{F}}^{2}.$ (7)
Here, $\mathrm{GL}(r,\mathbb{C})$ denotes the general linear group of degree $r$ over $\mathbb{C}$ and the invertible matrix ${\bm{Q}}$ is the best alignment between $({\bm{L}}_{k},{\bm{R}}_{k})$ and $(\bm{L}_{\star},\bm{R}_{\star})$. [36, Lemma 22] has shown that ${\bm{Q}}$ exists if $d_{k}$ is sufficiently small. In that case, the infimum in (7) can be replaced by the minimum, and the recovery error $\left\lVert{\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}\right\lVert_{\mathrm{F}}$ can be bounded by $d_{k}$ as stated in the following proposition.
Proposition 1.
For any $\varepsilon_{0}\in(0,1)$, provided $d_{k}<\varepsilon_{0}\sigma^{\star}_{r}$, then it holds that
@@
Proof.
The proof of this proposition is deferred to Section 5.2. ∎
The following Theorem 1 guarantees that our initialization (6) will start HSNLD sufficiently close to the ground truth, under some mild conditions. Thus, it is sufficient to bound $d_{k}$ for the recovery guarantee. In addition, our initial guess also satisfies the weak incoherence condition, which will be used and inherited iteratively in the convergence analysis. For ease of presentation, we denote the basin of attraction as: $\mathcal{E}(\delta,C):=\{({\bm{L}},{\bm{R}})~{}|~{}\mathrm{dist}({\bm{L}},{\bm{R}};\bm{L}_{\star},\bm{R}_{\star})\leq\delta~{}~{}\textnormal{and}~{}~{}\max\{\left\lVert{\bm{L}}{\bm{R}}^{*}\right\lVert_{2,\infty},\left\lVert{\bm{R}}{\bm{L}}^{*}\right\lVert_{2,\infty}\}\leq C\}$.
Theorem 1.
For any $\varepsilon_{0}\in(0,1)$, let $C=c\sqrt{c_{s}\mu r/n}\,\sigma^{\star}_{1}$ with $c\geq 1+\varepsilon_{0}$, and suppose Assumptions 1, 2 and 3 hold with $m\geq c_{0}\varepsilon_{0}^{-2}c_{s}\mu\kappa^{2}r^{2}\log n$ and $\alpha\leq\varepsilon_{0}/(50c_{s}\mu r^{1.5}\kappa)$, where $c_{0}$ is some universal constant. Then, the initialization step (6) satisfies $({\bm{L}}_{0},{\bm{R}}_{0})\in\mathcal{E}(\varepsilon_{0}\sigma^{\star}_{r},C)$, with probability at least $1-2n^{-2}$. In the case of full observation, probability rises to $1$.
Proof.
The proof of this theorem is deferred to Section 5.3. ∎
@@ 3.1 Full observation
We start the analysis with the simpler full observation case, i.e., $\Omega=[n]$. Theorem 2 presents the local linear convergence of HSNLD.
Theorem 2.
Suppose Assumptions 2 and 3 hold with $\alpha\lesssim\mathcal{O}(1/(c_{s}\mu r\kappa^{2}))$. Choose parameters $C=c\sqrt{c_{s}\mu r/n}\sigma^{\star}_{1}$ with $c\geq 1.01$, $\gamma_{k}\in[1+1/b_{0},2]$ with some fixed $b_{0}\geq 1$, and $\eta\in(0,1]$. Under full observation, for any $({\bm{L}}_{k},{\bm{R}}_{k})\in\mathcal{E}(0.01\sigma^{\star}_{r},C)$, it holds that
@@
Proof.
The proof of this theorem is deferred to Section 5.4.1. ∎
Combining Propositions 1, 1 and 2, we have the recovery guarantee, with a linear convergence, for the proposed HSNLD under full observation.
Corollary 1 (Recovery guarantee with full observation).
Suppose Assumptions 2 and 3 hold with $\alpha\lesssim\mathcal{O}(1/(c_{s}\mu r\kappa^{2}))$. Choose parameters $C=c\sqrt{c_{s}\mu r/n}\sigma^{\star}_{1}$ with $c\geq 1.01$, $\gamma_{k}\in[1+1/b_{0},2]$ with some fixed $b_{0}\geq 1$, and $\eta\in(0,1]$. Then, under full observation, in $K=\mathcal{O}(\log(1/\varepsilon))$ iterations, the outputs of Algorithm 1 satisfy
@@ 3.2 Partial observation
Now we will process the partial observation cases. One major challenge is the dependency among the entries on the antidiagonals of the Hankel matrix, which causes difficulty in uniformly bounding the subsampling error (see Lemma 12). Now, we present the local convergence under partial observation.
Theorem 3.
Suppose Assumptions 1, 2 and 3 hold with $m\gtrsim\mathcal{O}(c_{s}^{2}\mu^{2}r^{2}\kappa^{4}\log n)$ and $\alpha\lesssim\mathcal{O}(1/(c_{s}\mu r\kappa^{2}))$. Choose the parameters $C=c\sqrt{c_{s}\mu r/n}\sigma^{\star}_{1}$ with $c\geq 1.01$, $\gamma_{k}\in[1+1/b_{0},2]$ with some fixed $b_{0}\geq 1$, and $\eta\in(0,0.6]$. For any fixed $({\bm{L}}_{k},{\bm{R}}_{k})\in\mathcal{E}(0.01\sigma^{\star}_{r},C)$, with probability at least $1-8n^{-2}$ the following results hold $d_{k+1}\leq(1-0.6\eta)d_{k}~{}\textnormal{and}~{}({\bm{L}}_{k+1},{\bm{R}}_{k+1})\in\mathcal{E}(0.01\sigma^{\star}_{r},C)$.
Proof.
The proof of this theorem is deferred to Section 5.4.2. ∎
Combining Propositions 1, 1 and 3, we have the recovery guarantee, with a linear convergence, for the proposed HSNLD under partial observation.
Corollary 2 (Recovery guarantee with partial observation).
Suppose Assumptions 1, 2 and 3 hold with $m\gtrsim\mathcal{O}(c_{s}^{2}\mu^{2}r^{2}\kappa^{4}\log n)$ and $\alpha\lesssim\mathcal{O}(1/(c_{s}\mu r\kappa^{2}))$. Let the observation set $\Omega$ be resampled independently every iteration. Choose parameters $C=c\sqrt{c_{s}\mu r/n}\sigma^{\star}_{1}$ with $c\geq 1.01$, $\gamma_{k}\in[1+1/b_{0},2]$ with some fixed $b_{0}\geq 1$, and $\eta\in(0,0.6]$. Then, in $K=\mathcal{O}(\log(1/\varepsilon))$ iterations, the outputs of Algorithm 1 satisfy $\left\lVert{\bm{L}}_{K}{\bm{R}}_{K}^{*}-\mathcal{G}\bm{z}_{\star}\right\lVert_{\mathrm{F}}\leq 0.02(1-0.6\eta)^{K}\sigma^{\star}_{r}\leq\varepsilon$ with probability at least $1-\mathcal{O}(n^{-1})$, provided $\log(1/\varepsilon)\ll n$.
Remark 2.
Compared to the state-of-the-art HSGD which runs $\mathcal{O}(\kappa\log(1/\varepsilon))$ iterations for a $\varepsilon$-optimizer, the convergence rate of HSNLD is independent of the condition number $\kappa$. This makes HSNLD a powerful tool for heavily ill-conditioned Hankel applications, e.g., spectrally sparse signals with inseparable frequencies [30].
@@ 4 Numerical experiments
In this section, we will empirically demonstrate the performance of the proposed HSNLD with synthetic and real datasets. We compare our algorithm with HSGD [4], SAP [41], and RobustEMaC [19]. Parameters are set according to the original paper and further hand-tuned for best performance when applicable. Particularly, both HSNLD and HSGD use the same outlier sparsification parameters: $\gamma_{k}=1.05+0.45\times 0.95^{k}$ as suggested in [4], and the step size $\eta=0.5$. Moreover, we estimate the incoherence parameter $\mu$ via one-step Cadzow method [1] for all algorithms, and set the projection parameter $C$ according to Theorem 2 for HSNLD. The resampling is not implemented for HSNLD as it is just a proof technique and is not needed in practice. All results were conducted on Matlab with Intel i9-12950HX and 64GB RAM. The code is available at https://github.com/caesarcai/HSNLD.
@@ 4.1 Synthetic dataset
We generate an ill-conditioned Hankal matrix in its vector form $\bm{z}\in\mathbb{C}^{n}$ with $z_{t}=\sqrt{\varsigma_{t}}\sum^{r}_{j=1}a_{j}e^{2\pi\imath f_{j}t},$ where $\imath$ denotes the imaginary unit, $\{f_{j}\}\in\mathbb{R}$ is a randomly selected set of normalized off-grid frequencies, and $\{a_{j}\}\in\mathbb{R}$ is a set of $r$ evenly spaced numbers over $[1/\kappa,1]$. Thus, $\mathcal{G}\bm{z}\in\mathbb{C}^{n_{1}\times n_{2}}$ is exactly rank-$r$ with condition number $\kappa$. Then, $m$ observations are sampled uniformly without replacement from $\bm{z}$, and therein $\alpha m$ entries are randomly chosen for corruption. We add complex outliers to the corrupted entries whose real and imaginary parts are drawn uniformly over the intervals $[-10\mathbb{E}(|\mathrm{Re}(z_{t})|),10\mathbb{E}(|\mathrm{Re}(z_{t})|)]$ and $[-10\mathbb{E}(|\mathrm{Im}(z_{t})|),10\mathbb{E}(|\mathrm{Im}(z_{t})|)]$.
Phase transition. We demonstrate the recoverability and robustness of HSNLD via empirical phase transition under various settings. When an algorithm stops at $\|\mathcal{G}\Pi_{\Omega}\bm{z}_{k}+\mathcal{G}\Pi_{\Omega}\bm{s}_{k}-\mathcal{G}\Pi_{\Omega}\bm{f}\|_{\mathrm{F}}/\|\mathcal{G}\Pi_{\Omega}\bm{f}\|_{\mathrm{F}}\leq 10^{-5}$, we considered it a successful recovery if $\|\mathcal{G}\bm{z}_{k}-\mathcal{G}\bm{z}_{\star}\|_{\mathrm{F}}/\|\mathcal{G}\bm{z}_{\star}\|_{\mathrm{F}}\leq 10^{-3}$. For every problem setting, each algorithm runs for $20$ trials and reports the result as a pixel on the phase transition, where a white pixel means all $20$ trials were recovered and a black pixel means all $20$ trials were failed. The phase transition results are reported in Figures 3, 5 and 4 where one can see HSNLD delivers the best recoverability and robustness for ill-conditioned problems. Note that we only run the phase transition with a reasonably large $\kappa=10$ since the very ill-conditioned cases, such as $\kappa=2,000$, will take HSGD a forbidding long runtime.
@@
Convergence performance. In Figure 6, we plot the relative error v.s. runtime for the tested algorithms with different condition numbers. All problems have dimension $n=2^{16}-1$ and the results are averaged over 5 trials. When the condition number is perfect, i.e., $\kappa=1$, the HSNLD has tied convergence performance with HSGD. When the condition number grows, HSGD’s performance drops dramatically while HSNLD takes no impact—this matches our theoretical results. Generally speaking, SAP’s convergence performance is not impacted by the condition number either; however, the truncated SVD it applies every iteration may fail to function with vast condition number. RobustEMaC is not tested here because its sublinear convergence is too slow for the comparison.
@@
4.2 Direction of arrival estimation
Consider $r$ narrowband far-field source signals $\{\tilde{g}_{i}\}_{i=1}^{r}\subseteq\mathbb{C}$ impinging on an array of omnidirectional sensors from distinct yet close directions as shown in Figure 1. Let the DOAs corresponding to these sources be denoted by $\{\theta_{i}\}_{i=1}^{r}$. The noiseless received signal $x_{j}$ at the $j$-th sensor of a Uniform Linear Array, with a half-wavelength spacing between adjacent sensors, is given by:
$x_{j}=\sum_{i=1}^{r}\tilde{g}_{i}e^{-\pi\imath(j-1)\sin\theta_{i}},\quad j=1,2,\dots,n,\vspace{-0.05in}$
where $\imath=\sqrt{-1}$. Consider the scenario where only a subset $\Omega\subseteq[n]$ of the sensors is available, and the received signals are partially corrupted. The resulting observations are of the form $y_{j}=x_{j}+o_{j},~{}j\in\Omega$, where $o_{j}$ represents the outlier at the $j$-th sensor. The problem of recovering $\bm{x}=[x_{1},x_{2},\cdots,x_{n}]^{\top}\in\mathbb{C}^{n}$ from observed data on $\Omega$ (i.e., $\{y_{j}\}_{j\in\Omega}$) is equivalent to recover $\bm{z}=\mathcal{W}^{-1}\bm{x}$ as defined in eq. 1 and can be formulated as a robust rank-$r$ Hankel matrix completion problem. For the interested reader, please see [39, Section 6.3] for further details.
We test HSNLD against the state-of-the-art with DOA data generated from $n=2^{12}$ sensors and $r=3$ far-field sources, with directions $\theta=87^{\circ},87.1^{\circ},87.3^{\circ}$ as shown in Figure 1. The underlying square Hankel matrix corresponding to the DOA data has a condition number $\kappa\approx 1083$. In this experiment, only $1.5\%$ sensors are activated (or deployed), and $10\%$ of the collected data are corrupted by complex outliers whose components are uniformly drawn from the intervals $[-\sum_{j=1}^{n}\mathrm{Re}(x_{j})/n,\sum_{j=1}^{n}\mathrm{Re}(x_{j})/n]$ and $[-\sum_{j=1}^{n}\mathrm{Im}(x_{j})/n,\sum_{j=1}^{n}\mathrm{Im}(x_{j})/n]$. All test algorithms are terminated at $\|z_{k}-z\|/\|z\|\leq 10^{-5}$. Figure 7 plots the signal recovered by HSNLD, compared to the original signal and corrupted partial observations. HSNLD successfully recovers the ground truth. The state-of-the-art algorithms, HSGD and SAP, also successfully recover the DOA signal. Due to space constraints, their plots are omitted. However, the runtime reported in Table 1 shows the significant speed advantage of HSNLD; it achieves nearly 200×200$\times$200 × speedup to HSGD and 10×10$\times$10 × speedup to SAP.
Figure 7: DOA data recover by HSNLD. Only $1.5\%$ of the sensors are activated/deployed and $10\%$ of the observations are corrupted by outliers. Left and right: Real and imaginary parts of the data.
Table 1: The number of iterations and runtime for DOA data recovery
Method HSNLD HSGD SAP
Iteration 72 29186 32
Runtime (sec) 0.236 38.611 2.040
4.3 Nuclear magnetic resonance
We test HSNLD with nuclear magnetic resonance (NMR) signal recovery problem, a classic application of low-rank Hankel recovery. The 1D NMR signal being tested has a length of $32,768$ with an approximate rank of $40$ and condition number $30$. The observation rate is set to be $30\%$ and $10\%$ outliers are added. All algorithms stop at $|\mathrm{err}_{k+1}-\mathrm{err}_{k}|/\mathrm{err}_{k}<10^{-4}$. The recovery results are reported in Figures 8 and 2. One can see that all three tested algorithms recovered the signal successfully while HSNLD is significantly faster.